# Latent Dataset Distillation with Diffusion Models

### v2

@@ 4.2 Results
Cross-Architecture Evaluation, IPC=10. Further supporting our findings, Table 2 illustrates the superior performance of LD3M over GLaD in scenarios with IPC=10 using the distillation algorithms DC and DM on ImageNet A to E. The trend observed is similar to the IPC=1 scenario, with LD3M demonstrating a clear advantage over GLaD. The exception of DM on ImageNet-C is marginal and still outperforms pixel-based distillation, but also shows that the application of LDM for IPC=10 has the most significant improvements in DM compared to DC, e.g., ca. 4 p.p. improvement on ImageNet-A. LD3M reaches an overall average accuracy of 47.08 % with DC and 47.28 % with DM compared to 45.88 % and 45.76 % of GLaD with StyleGAN-XL. This experiment highlights the consistency and reliability of the LDM approach across different settings.
High-Resolution Distillation. In the context of $256\times 256$ data, as shown in Table 3, a similar pattern of improved performance is evident. LD3M achieved all top-3 results per subset. To explore the versatility of LDMs, we experimented with various training configurations, including models pre-trained on ImageNet, FFHQ, and a randomly initialized model. Consistent with the findings using GANs, these varied LDM configurations demonstrated improved cross-architectural generalization compared to direct pixel-space distillation.
Figure 3: Example $256\times 256$ images of a distilled class (ImageNet-B: Lorikeet) with differently initialized GLaD and LD3M. Initializations are denoted at the bottom.
Visual Comparison. As presented in Figure 2, visual comparisons depict the results of GLaD and the LD3M. These comparisons reveal that the generated images by LD3M are notably more abstract and exhibit higher contrast than those produced by GLaD. While GLaD tends to generate images more closely resembling the initial real image used in the latent code representation, LD3M diverges towards a more stylized and vivid visual representation. This highlights a more flexible learning space for dataset distillation algorithms.
@@
Diffusion Time Steps Analysis. Our investigation into the trade-off between the runtime and accuracy of distilled datasets is shown in Figure 4. It illustrates the diffusion steps versus the accuracy of our LDM-based distillation method across different maximum time steps $T$ in the diffusion process, specifically evaluated on the ImageNet A-E datasets using the MTT algorithm. There is a notable increase in accuracy for setting the maximum time step up to 40 for the diffusion process. However, as we progress beyond 40 time steps, performance is significantly declining. An example is shown in Figure 5 for the freight car class from ImageNet-C. This decline could be attributed to vanishing gradients during the diffusion process, suggesting that strategies other than the linear addition of the initial state could avoid the performance decline.
However, if the same pre-trained LDM is applied, we do not expect other strategies to improve the overall reachable image quality significantly. The ability to tailor the diffusion process according to specific accuracy and runtime opens up new possibilities for dataset distillation across various computational settings. It allows for strategically allocating computational resources, ensuring the models are accurate and practically viable for real-world applications.
Figure 4: Distilled dataset performance (average of ImageNet A-E via MTT and IPC=1) for different max. time steps employed in the diffusion process with mean and standard deviation (light blue dotted). It illustrates the trade-off between runtime and accuracy over increasing time steps. LD3M shows improved performance for the first 40 time steps, while the performance deteriorates for greater time steps. GLaD needs around 140 seconds per iteration on the same hardware (NVIDIA RTX3090) with 34.72 % average accuracy, shown as dotted lines. Best trade-off marked with "X".
@@
Table 4: Comparison of different LDM variations with MTT on ImageNette for 1,000 distillation steps and IPC=1. Tested were the distillation of the LDM by making only the conditioning learnable, one LDM with also learnable initial latent representation ($\mathbf{z}_{T}$), and lastly, one LDM which incorporates the initial latent representation in the calculation of intermediate latent states (see Equation 6).
@@ Acknowledgment
Iteration (Distillation) 5,000 ($128\times 128$) / 10,000 ($256\times 256$
@@
Classes in Subsets
Table 5 list ImageNet classes that are used in the subsets employed in this work.
@@
Figure 6: Images distilled by MTT in LD3M for IPC=1.
Figure 7: Images distilled by DC in LD3M for IPC=1.
Figure 8: Images distilled by DM in LD3M for IPC=1.
Figure 9: Images distilled by DC in LD3M for IPC=10 and ImageNet-A.
Figure 10: Images distilled by DC in LD3M for IPC=10 and ImageNet-B.
Figure 11: Images distilled by DC in LD3M for IPC=10 and ImageNet-C.
Figure 12: Images distilled by DM in LD3M for IPC=10 and ImageNet-A.
Figure 13: Images distilled by DM in LD3M for IPC=10 and ImageNet-B.
Figure 14: Images distilled by DM in LD3M for IPC=10 and ImageNet-C.

### v3

@@ 4.2 Results
Dist. Method Dist. Space AlexNet ResNet18 VGG11 ViT Average
DC Pixel 25.9$\pm$0.2 27.3$\pm$0.5 28.0$\pm$0.5 22.9$\pm$0.3 26.0$\pm$0.4
GLaD (Rand G) 30.1$\pm$0.5 27.3$\pm$0.2 28.0$\pm$0.9 21.2$\pm$0.6 26.6$\pm$0.5
GLaD (Trained G) 26.0$\pm$0.7 27.6$\pm$0.6 28.2$\pm$0.6 23.4$\pm$0.2 26.3$\pm$0.5
LD3M (Trained G) 27.2$\pm$0.8 26.6$\pm$0.9 31.5$\pm$0.3 29.0$\pm$0.2 28.6$\pm$0.6
DM Pixel 22.9$\pm$0.2 22.2$\pm$0.7 23.8$\pm$0.5 21.3$\pm$0.5 22.6$\pm$0.5
GLaD (Rand G) 23.7$\pm$0.3 21.7$\pm$1.0 24.3$\pm$0.8 21.4$\pm$0.2 22.8$\pm$0.6
GLaD (Trained G) 25.1$\pm$0.5 22.5$\pm$0.7 24.8$\pm$0.8 23.0$\pm$0.1 23.8$\pm$0.5
LD3M (Trained G) 27.2$\pm$0.4 17.0$\pm$0.7 25.4$\pm$0.4 23.8$\pm$0.3 23.4$\pm$0.5
Table 4: CIFAR-10 Performance on unseen architectures, IPC=1.
@@
Cross-Architecture Evaluation, IPC=10. Further supporting our findings, Table 2 illustrates the superior performance of LD3M over GLaD in scenarios with IPC=10 using the distillation algorithms DC and DM on ImageNet A to E. The trend observed is similar to the IPC=1 scenario, with LD3M demonstrating a clear advantage over GLaD. The exception of DM on ImageNet-C is marginal and still outperforms pixel-based distillation, but also shows that the application of LDM for IPC=10 has the most significant improvements in DM compared to DC, e.g., ca. 4 p.p. improvement on ImageNet-A. LD3M reaches an overall average accuracy of 47.08 % with DC and 47.28 % with DM compared to 45.88 % and 45.76 % of GLaD. This experiment highlights the consistency and reliability of the LDM approach across different settings.
High-Resolution. In the context of $256\times 256$ data, as shown in Table 3, a similar pattern of improved performance is evident. LD3M achieved all top-3 results per subset. To explore the versatility of the LDMs that can be used in LD3M, we experimented with various training configurations, including models pre-trained on ImageNet, FFHQ, and a randomly initialized model (no pre-training). Consistent with the findings using GANs, these varied LDM configurations demonstrated improved cross-architectural generalization compared to direct pixel-space distillation.
CIFAR-10. Like Cazenavette et al. [8], we evaluated LD3M on CIFAR-10 in Table 4. While it presents a less challenging task (lower resolution), we observe that LD3M is still on par with DM while significantly better with DC.
Initializations. In Table 5, we investigate the influence of different initializations of the latent and conditioning code, which was either Gaussian noise or initialized with a randomly selected image and its label of the respective class. As a result, we can see that the performance of LD3M benefits from real images, especially for the distillation algorithms DM and DC, while MTT seems more robust.
Dist. Method Dist. Space AlexNet ResNet18 VGG11 ViT Average
MTT Gauss. noise 28.7$\pm$1.6 34.1$\pm$1.5 32.2$\pm$0.6 29.1$\pm$1.8 31.0$\pm$1.4
Random image 30.1$\pm$1.4 35.6$\pm$1.4 32.2$\pm$0.3 30.0$\pm$1.2 32.0$\pm$1.3
DC Gauss. noise 13.1$\pm$1.5 11.6$\pm$1.8 13.8$\pm$2.2 13.7$\pm$2.7 13.1$\pm$2.1
Random image 31.6$\pm$1.3 30.4$\pm$0.6 31.8$\pm$1.2 37.7$\pm$1.5 32.9$\pm$2.1
DM Gauss. noise 13.4$\pm$2.0 12.4$\pm$1.8 13.4$\pm$1.4 14.4$\pm$2.1 13.4$\pm$1.8
Random image 31.9$\pm$1.3 23.2$\pm$2.2 25.9$\pm$2.0 26.1$\pm$1.4 26.8$\pm$1.7
Table 5: Performance with different initializations of the latent and conditioning codes per class on ImageNette for the IPC=1 setting with 5,000 iterations. MTT with random initialization reaches comparable performance (still under initialization with randomly selected images and their labels), whereas random initialization leads to synthetic datasets with significantly reduced quality for DC and DM.
Figure 3: Example $256\times 256$ images of a distilled class (ImageNet-B: Lorikeet) with differently initialized generators GLaD and LD3M. The various initializations, i.e., which dataset was used for training the generators, are denoted at the bottom.
Visual Comparison. As presented in Figure 2, visual comparisons depict the results of GLaD and the LD3M. These comparisons reveal that the generated images by LD3M are notably more abstract and exhibit higher contrast than those produced by GLaD. This highlights a more flexible learning space for dataset distillation algorithms.
@@
Diffusion Steps Analysis. Our investigation on the trade-off between runtime and accuracy is shown in Figure 4. It illustrates the diffusion steps versus the accuracy of our LDM-based distillation method across different maximum time steps $T$ in the diffusion process, specifically evaluated on the ImageNet A-E datasets using the MTT algorithm. There is a notable increase in accuracy for setting the maximum time step up to 40 for the diffusion process. However, performance is significantly declining as we progress beyond 40. An example is shown in Figure 5 for the freight car class from ImageNet-C. This decline could be attributed to vanishing gradients during the diffusion process, suggesting that strategies other than the linear addition of the initial state could avoid the performance decline.
Overall, LD3M with $T=20$ needs 574 minutes versus 693 minutes for GLaD. Furthermore, on an A100-40GB, LD3M with $T=20$ needs 73.58% max. GPU allocation / 29.4GB vs GLaD with 77.94% max. GPU allocation / 31.2GB. Also, LD3M can distill with less memory by adjusting the number of diffusion steps (i.e., from $T=2$ to $T=35$ ), which is not doable with GLaD (fixed GAN). The ability to tailor the diffusion process according to specific accuracy and runtime opens up new possibilities for dataset distillation across various computational settings. It allows for strategically allocating computational resources, ensuring the models are accurate and practically viable for real-world applications. While dataset distillation remains computationally expensive, LD3M introduces a flexible and significantly lighter alternative to GLaD.
Figure 4: Distilled dataset performance (average of ImageNet A-E via MTT and IPC=1) for different max. time steps employed in the diffusion process with mean and standard deviation (light blue dotted). It illustrates the trade-off between runtime and accuracy over increasing time steps. LD3M shows improved performance for the first 40 time steps, while the performance deteriorates for greater time steps. GLaD needs around 140 seconds per iteration on the same hardware (NVIDIA RTX3090) with 34.72 % average accuracy, shown as dotted lines. Best trade-off marked with ”X”.
@@
Table 6: Comparison of different LDM variations with MTT on ImageNette for 1,000 distillation steps and IPC=1. Tested were the distillation of the LDM by making only the conditioning learnable, one LDM with also learnable initial latent representation ($\mathbf{z}_{T}$), and lastly, one LDM which incorporates the initial latent representation in the calculation of intermediate latent states.
@@ 6 Limitations
7 Societal Impact
We focus on publishing a new generative model for dataset distillation. As such, the generation of synthetic images and their associated privacy benefits pose no immediate threat to individuals or organizations. Yet, the application of dataset distillation might have an environmental impact (e.g., using huge models for image distillation).
@@ Acknowledgment
Iteration (Distillation) 5,000 ($128\times 128$) / 10,000 ($256\times 256$)
@@
Table 7: Common hyper-parameters for training the distillation algorithms used in this work.
@@
Figure 6: Images distilled by MTT in LD3M for IPC=1.
Figure 7: Images distilled by DC in LD3M for IPC=1.
Figure 8: Images distilled by DM in LD3M for IPC=1.
Figure 9: Images distilled by DC in LD3M for IPC=10 and ImageNet-A.
Figure 10: Images distilled by DC in LD3M for IPC=10 and ImageNet-B.
Figure 11: Images distilled by DC in LD3M for IPC=10 and ImageNet-C.
Figure 12: Images distilled by DC in LD3M for IPC=10 and ImageNet-D.
Figure 13: Images distilled by DC in LD3M for IPC=10 and ImageNet-E.
Figure 14: Images distilled by DM in LD3M for IPC=10 and ImageNet-A.
Figure 15: Images distilled by DM in LD3M for IPC=10 and ImageNet-B.
Figure 16: Images distilled by DM in LD3M for IPC=10 and ImageNet-C.
Figure 17: Images distilled by DM in LD3M for IPC=10 and ImageNet-D.
Figure 18: Images distilled by DM in LD3M for IPC=10 and ImageNet-E.